OpenStudy (anonymous):

A wooden garden tray with a base area of xy square inches is filled with soil up to a certain height. The tray contains (x2y + 10xy2 + 12xy) cubic inches of soil. The height of the tray is represented by the expression x squared y plus 10 x y squared plus 12 x y all over xy Which expression below is the correct simplified form of the height of the tray? 11xy + 12 x + 10y + 12 x + 10xy2 + 12xy x2y + 10xy2 + 12

OpenStudy (skullpatrol):

The tray contains $$x^2y + 10xy^2 + 12xy$$ cubic inches of soil. The height of the tray is represented by the expression $$\frac{x^2 y + 10 x y^2 + 12 x y}{ xy}$$ Which expression below is the correct simplified form of the height of the tray? $$11xy + 12$$ $$x + 10y + 12$$ $$x + 10xy^2 + 12xy$$ $$x^2y + 10xy^2 + 12$$

OpenStudy (anonymous):

the last one

OpenStudy (mertsj):

\[\frac{x^2y+10xy^2+12xy}{xy}=\frac{x^2y}{xy}+\frac{10xy^2}{xy}+\frac{12xy}{xy}\]